OpenStudy (anonymous):

simplify 3x2-3x-6 over x2-4

OpenStudy (anonymous):

\[\frac{3x^2-3x-6}{x^2-4}=\frac{3(x-2)(x+1)}{(x+2)(x-2)}=\frac{3(x+1)}{x+2}\]\]